An item for sale is marked down 20%. By what percent must it then be marked
up in order to return to the original selling price?

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

An item for sale is marked down 20%. By what percent must it then be marked
up in order to return to the original selling price?
.
Let x=original price
then the sale price is
x - .20x
x(1-.20)
x(.80)
.80x
.
Let y = mark up to get back to the original price
then
.80x * y = x
y = x/(.80x)
y = 1/.80
y = 1.25
or
y = 125%

Answer by Alan3354(69443)

It must be marked up 25%, not 125%.
-------------------
x*(0.80)*(1.25) = x